OpenStudy (anonymous):

EASY FAN/MEDAL Solve for x: |2x + 6| − 4 = 20 A)x = 9 and x = 11 B)x = −9 and x = 15 C)x = 9 and x = −15 D)No solution

OpenStudy (anonymous):

If my memory is correct, you would need to split this up into two problems: \[\left| 2x + 6 \right| -4 = 20\] and \[\left| 2x - 6\right|+4 =20\] Then solve for x for each equation...

OpenStudy (anonymous):

fan and medal for the answer

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the answer is A

OpenStudy (anonymous):

just do this on your calculator and you'll see... :)

OpenStudy (anonymous):

thank you sooo much ^_^

OpenStudy (anonymous):

np :)

myininaya (myininaya):

\[|2x-6|-4=20 \\ \text{ add 4 on both sides } \\ |2x+6|=24 \\ \text{ this gives two equations to solve } \\ 2x+6=24 \text{ or } 2x+6=-24 \]